# Simulating Molecular Single Vibronic Level Fluorescence Spectra with ab initio Hagedorn Wavepacket Dynamics

### v4

@@ Abstract
We present a practical, ab initio time-dependent method using Hagedorn wavepackets to efficiently simulate single vibronic level (SVL) fluorescence spectra of polyatomic molecules from arbitrary initial vibrational levels. We apply the method to compute SVL spectra of anthracene by performing wavepacket dynamics on a 66-dimensional harmonic potential energy surface constructed from density functional theory calculations. The Hagedorn approach captures both mode distortion and Duschinsky rotation within the harmonic approximation. We not only reproduce the previously reported simulation results for singly excited $12^{1}$ and $\overline{11}^{1}$ levels, but are also able to compute SVL spectra from multiply excited levels in good agreement with experiments. Notably, all spectra were obtained from the same wavepacket trajectory without any additional propagation beyond what is required for the emission spectrum from the ground vibrational level of the electronically excited state.
@@ 1 Introduction
Calculations of SVL spectra have mostly relied on the time-independent approach, where individual Franck–Condon factors are computed for each vibronic transition.10, 11, 17, 16 However, for larger polyatomic molecules, the task becomes challenging due to the large number of possible transitions and the computation of nonseparable overlap integrals.18, 19, 20, 21 The time-dependent approach, which is more natural for low- or intermediate-resolution spectra, avoids the need to prescreen transitions and can more efficiently accommodate Duschinsky rotation (mode mixing),22, 23, 14 anharmonicity,24, 25, 26, 27 Herzberg-Teller,28, 29, 22, 30, 31, 25, 26, 32, 33 and temperature effects34, 35, 36, 37 in both linear and multidimensional38, 39, 40, 41 vibronic spectra.
@@
In order to evaluate SVL spectra associated with emission from arbitrary vibronic levels, we recently proposed42 another time-dependent approach, one based on Hagedorn functions.43, 44, 45 These functions in the form of a Gaussian multiplied by carefully constructed polynomials46, 47, 45 are exact solutions of the time-dependent Schrödinger equation with a quadratic potential and have drawn attention in applied mathematics48, 49, 44, 50, 51, 52, 46, 53, 54, 45 because of their promising applications in physics and chemistry.55, 56, 57, 58, 59
The recursive expressions we derived for their overlap integrals60 made it possible to apply Hagedorn wavepackets in computational spectroscopy.42, 61 While Hagedorn functions, as a complete orthonormal basis, can represent initial states of arbitrary shape, their form is particularly well-suited for simulating SVL spectra. Within the harmonic and Condon approximations, the SVL process may be represented by a single Hagedorn function at all times and provides a clear, uncomplicated first demonstration of the Hagedorn dynamics approach to vibronic spectroscopy. In ref 42, we validated the Hagedorn approach to compute SVL spectra against quantum split-operator results in two-dimensional harmonic model systems incorporating displacement, mode distortion, and Duschinsky rotation effects.
Here, we describe how the approach from ref 42 extends to realistic polyatomic molecules within the global harmonic approximation, which provides a convenient starting point for studying larger molecular systems and can guide the selection of ab initio methods for further on-the-fly local harmonic studies that include anharmonic effects.61 In order to implement ab initio Hagedorn wavepacket dynamics, we neglect ro-vibrational coupling and perform the vibrational dynamics in the normal-mode coordinates on a full-dimensional, global harmonic potential energy surface derived from electronic structure calculations. As a test, we simulate the SVL spectra of anthracene from $\overline{11}^{j}$ and $12^{j}$ ($j=1,2$) levels, for which assigned experimental data are available.3 We also predict the SVL spectra from the $12^{1}6^{1}$, $12^{1}\overline{5}^{1}$, and $12^{1}5^{1}$ levels, which remain to be measured experimentally. Whereas the singly excited cases were already successfully simulated by Tapavizca,14 our Hagedorn approach can treat higher and mixed excitations at a negligible additional cost.
@@ 2.1 Time-dependent approach to spectroscopy
In the time-dependent approach to vibronic spectroscopy, the spectrum is evaluated as the Fourier transform of an appropriate wavepacket autocorrelation function
@@
[…]
@@ 2.2 Representation of SVL initial states by Hagedorn functions
in Hagedorn’s parametrization, where $x:=q-q_{t}$ is the shifted position, $q_{t}$ and $p_{t}$ are the position and momentum of the center of the wavepacket, $S_{t}$ is the classical action, and $Q_{t}$ and $P_{t}$ are complex-valued $D$-dimensional matrices that satisfy the so-called symplecticity conditions52, 45 (see eqs 5 and 6 of ref 42) and determine the width matrix $P\cdot Q^{-1}$ of the Gaussian. We can then apply Hagedorn’s raising operator
@@
in the form of multivariate polynomials multiplied by a common Gaussian, with $\langle j\rangle=(\underbrace{0,\dots,0}_{j-1},1,\underbrace{0,\dots,0}_{D-j})$ being the $D$-dimensional unit vector along the $j$-th degree of freedom.62, 43, 52, 45
The initial vibrational state $|\psi_{0}\rangle=|{K}\rangle$ in the SVL emission process can be exactly represented by a single Hagedorn function $\varphi_{K}$ in the normal-mode coordinates. In the harmonic approximation, normal-mode coordinates for a given electronic state diagonalize the Hessian at its equilibrium geometry, and the ground vibrational wavefunction is a Gaussian (eq 3) with a diagonal width matrix $P_{0}\cdot Q_{0}^{-1}$. With both $Q_{0}$ and $P_{0}$ diagonal, the associated multidimensional Hagedorn function is a product of univariate functions, each of which is a Gaussian multiplied by a Hermite polynomial. As we will see in the next sections, this simple Hermite factorization is lost during the dynamics on a general harmonic surface.
@@ 2.3 Hagedorn wavepacket dynamics
Remarkably, Hagedorn wavepackets are exact solutions to the time-dependent Schrödinger equation with a harmonic potential
@@
$\dot{Q_{t}}$ $=m^{-1}\cdot P_{t},$ $\dot{P_{t}}$ $=-\kappa_{g}\cdot Q_{t},$
$\dot{S_{t}}$ $=L_{t}=p^{T}\cdot m^{-1/2}\cdot p/2-V_{g}(q_{t}),$ (7)
where $m$ is the mass matrix and $L_{t}$ is the Lagrangian.52, 45, 63 These equations of motion are the same as in the thawed Gaussian approximation,64, 65 which has been applied to vibronic spectra arising from ground vibrational levels23, 24, 66, 67 and can be extended to include a linear polynomial prefactor.68, 30, 35, 32, 69, 70 Using Hagedorn wavepackets, it is possible to treat arbitrary polynomials times a Gaussian (describing, e.g., excited vibrational states), and since the propagation (7) still depends only on the Gaussian’s parameters, a single Gaussian trajectory is sufficient to obtain SVL spectra from all vibrational levels.
@@ 2.4 Combination with ab initio electronic structure calculations
The first step in transforming Cartesian coordinates to mass-weighted normal-mode coordinates is minimizing the ro-vibrational coupling relative to a reference geometry. This is achieved by moving to the center-of-mass frame and applying the Kabsch algorithm to satisfy the translational and rotational Eckart conditions.71, 72, 73 Next, the normal modes and the corresponding transformation matrix are found by diagonalizing the mass-weighted Hessian matrix of the reference state and projecting out the rotational and translational degrees of freedom (see section 6.7 of ref 74). In the SVL application, the reference geometry and Hessian are taken from the initial, excited electronic state so that each vibrational eigenstate $|K\rangle$ may be represented by a single Hagedorn function $\varphi_{K}$. The diagonalized vibrational Hessian $\kappa_{e}$ in the excited-state normal-mode coordinates determines the width matrix $\mathrm{A_{0}}=i\sqrt{m^{1/2}\cdot\kappa_{e}\cdot m^{1/2}}$ of the Gaussian $\varphi_{0}$ associated with the SVL initial states. In Hagedorn’s parametrization, we set $Q_{0}=(\operatorname{Im}\mathrm{A}_{0})^{-1/2}$ and $P_{0}=\mathrm{A}_{0}\cdot Q_{0}$ to satisfy the symplecticity relations.
In the adiabatic harmonic approximation, the Hessian $\kappa_{g}$ defining the ground-state potential (eq 6) used for propagation is evaluated at the ground-state equillibrium geometry $q_{\text{ref},g}=q_{\text{eq},g}$; both the geometry and Hessian are transformed to be in the excited-state mass-weighed normal-mode coordinates. In the vertical harmonic approximation, the ground-state Hessian $\kappa_{g}$ is instead evaluated at the optimized excited-state geometry $q_{\text{ref},g}=q_{\text{eq},e}$. Due to mode mixing, $\kappa_{g}$ is in general not diagonal when expressed in the excited-state normal-mode coordinates. The initially diagonal $Q_{t=0}$ matrix of the SVL initial state will then no longer be diagonal after being evolved with the ground-state potential, and the propagated Hagedorn function will not remain a simple product of univariate Hermite functions.46, 47, 45
@@ 3 Computational details
All geometry optimization and frequency calculations were performed using Gaussian 1675 at the same level of theory (PBE0/def2-TZVP76, 77) as in ref 14 for the ground ($S_{0}\,\mathrm{{}^{1}A_{g}}$) and the first excited ($S_{1}\,\mathrm{{}^{1}B_{2u}}$) electronic states of anthracene. The optimized structures of the two states and the frequencies of the vibrational modes analyzed in this work are listed in the Supporting Information. The assignment of the vibrational modes follows the convention of ref 3 and the Supporting Information of ref 14.
Although the equations of motion (7) can also be solved analytically to obtain a solution at an arbitrarily long time $t$ in a harmonic system,78 we used a second-order TVT geometric integrator 79, 52, 63 to propagate the parameters of the Gaussian for a total time of $8\times 10^{5}\,\text{au}$ ($\sim$19 ps) with a time step of 8 au in the global harmonic systems considered in this work. Since the spectrum (2) depends on the values of the autocorrelation function $C(t)$ at all times $t$ from 0 to $t_{\text{max}}$ (the time that gives the desired spectral resolution), the numerical approach is nearly as accurate and efficient as the analytical one and, moreover, offers the flexibility for future treatment of anharmonic effects.
After the (single!) trajectory of the parameters of the Gaussian was obtained in a given harmonic system, the autocorrelation functions between Hagedorn wavepackets were computed every four steps for each initial vibrational level ($0^{0},12^{1},12^{2},\overline{11}^{1},\overline{11}^{2},12^{1}6^{1},12^{1}\overline{5}^{1}$, and $12^{1}5^{1}$) using the algebraic algorithm described in ref 60. Given the limited range of the observed experimental data, it was unnecessary to extend the range of the simulated spectra by computing the autocorrelation function at every step. A Gaussian damping function with a half-width at half-maximum of 20000 au was applied to the autocorrelation functions before performing the Fourier transform.
To facilitate the comparison, the intensities of the highest peaks in all spectra (simulated and experimental) were set to unity. A horizontal shift, determined by aligning the $0^{0}_{0}$ peak in the computed ground-level ($0^{0}$) fluorescence spectrum (see the Supporting Information) to be at the experimental origin $27709\,\text{cm}^{-1}$, was applied to the wavenumbers of the simulated spectra in global harmonic potentials to correct for the relatively large error (compared to the vibrational features) in the computed electronic excitation energy; the $\omega^{3}$ factor in eq 2 was then adjusted based on the corrected wavenumbers (see the Supplemental Material of ref 61 for details).

### v5

@@ Abstract
We present a practical, ab initio time-dependent method using Hagedorn wavepackets to efficiently simulate single vibronic level (SVL) fluorescence spectra of polyatomic molecules from arbitrary initial vibrational levels. We apply the method to compute SVL spectra of anthracene by performing wavepacket dynamics on a 66-dimensional harmonic potential energy surface constructed from density functional theory calculations. The Hagedorn approach captures both mode distortion (frequency changes) and mode mixing (Duschinsky rotation) within the harmonic approximation. We not only reproduce the previously reported simulation results for singly excited $12^{1}$ and $\overline{11}^{1}$ levels, but are also able to compute SVL spectra from multiply excited levels in good agreement with experiments. Notably, all spectra were obtained from the same wavepacket trajectory without any additional propagation beyond what is required for the emission spectrum from the ground vibrational level of the electronically excited state.
@@ 1 Introduction
Calculations of SVL spectra have mostly relied on the time-independent approach, where individual Franck–Condon factors are computed for each vibronic transition.10, 11, 17, 16 However, for larger polyatomic molecules, the task becomes challenging due to the large number of possible transitions and the computation of nonseparable overlap integrals.18, 19, 20, 21 The time-dependent approach, which is more natural for low- or intermediate-resolution spectra, avoids the need to prescreen transitions and can more efficiently accommodate Duschinsky rotation (mode mixing),22, 23, 14 anharmonicity,24, 25, 26, 27 Herzberg-Teller,28, 29, 22, 30, 31, 25, 26, 32, 33 and temperature effects34, 35, 36, 37, 38 in both linear and multidimensional39, 40, 41, 42 vibronic spectra.
@@
In order to evaluate SVL spectra associated with emission from arbitrary vibronic levels, we recently proposed43 another time-dependent approach, one based on Hagedorn functions.44, 45, 46 These functions in the form of a Gaussian multiplied by carefully constructed polynomials47, 48, 46 are exact solutions of the time-dependent Schrödinger equation with a quadratic potential and have drawn attention in applied mathematics49, 50, 45, 51, 52, 53, 47, 54, 55, 46 because of their promising applications in physics and chemistry.56, 57, 58, 59, 60
The recursive expressions we derived for their overlap integrals61 made it possible to apply Hagedorn wavepackets in computational spectroscopy.43, 62 While Hagedorn functions, as a complete orthonormal basis, can represent initial states of arbitrary shape, their form is particularly well-suited for simulating SVL spectra. Within the harmonic and Condon approximations, the SVL process may be represented by a single Hagedorn function at all times and provides a clear, uncomplicated first demonstration of the Hagedorn dynamics approach to vibronic spectroscopy. In ref 43, we validated the Hagedorn approach to compute SVL spectra against quantum split-operator results in two-dimensional harmonic model systems incorporating displacement, mode distortion (changes of vibrational frequencies in different electronic states), and mode mixing (Duschinsky rotation) effects.
Here, we describe how the approach from ref 43 extends to realistic polyatomic molecules within the global harmonic approximation, which provides a convenient starting point for studying larger molecular systems and can guide the selection of ab initio methods for further on-the-fly local harmonic studies that include anharmonic effects.62 In order to implement ab initio Hagedorn wavepacket dynamics, we neglect ro-vibrational coupling and perform the vibrational dynamics in the normal-mode coordinates on a full-dimensional, global harmonic potential energy surface derived from electronic structure calculations. As a test, we simulate the SVL spectra of anthracene from $\overline{11}^{j}$ and $12^{j}$ ($j=1,2$) levels, for which assigned experimental data are available.3 We also predict the SVL spectra from the $12^{1}6^{1}$, $12^{1}\overline{5}^{1}$, and $12^{1}5^{1}$ levels, which remain to be measured experimentally. Whereas the singly excited cases were already successfully simulated by Tapavizca,14 our Hagedorn approach can treat higher and mixed excitations from a single semiclassical trajectory by computing overlaps between Hagedorn functions at a small additional cost that becomes negligible in on-the-fly local harmonic applications.
@@ 2.1 Time-dependent approach to spectroscopy
The traditional “sum-over-states” approach for evaluating vibronic spectra relies on computing the Franck–Condon overlaps between the initial and final vibrational states. In displaced and distorted harmonic potentials, these overlap integrals can be evaluated exactly using analytical formulas,63, 64 or, when Duschinsky effects are considered, through recursive procedures.65, 66, 67 While this time-independent approach provides an explicit contribution from each vibronic transition to the spectrum, computational cost becomes prohibitive in larger molecules if one does not restrict the number of final vibrational states included in the calculation (either heuristically or through an automated procedure implemented in packages such as FCclasses 66, 67).
An alternative, time-dependent approach avoids the need to pre-select and compute individual transitions. The spectrum is instead evaluated as the Fourier transform of an appropriate wavepacket autocorrelation function
@@
$\sigma_{\text{em}}(\omega)=\frac{4\omega^{3}}{3\pi\hbar c^{3}}|{\mu}_{ge}|^{2}\operatorname{Re}\int^{\infty}_{0}\overline{C(t)}\exp[it(\omega-\omega_{e,K})]\,dt,$ (2)
@@
In contrast to the time-independent approach, the dynamics-based approach does not explicitly enumerate the final vibrational states in the vibronic spectrum and obtains all peaks at once. The convergence of the spectrum (in terms of the frequency range and resolution) is determined by the time step used for evaluating autocorrelation functions and the total time of propagation. Although the individual contributions of each transition are not directly available, the time-dependent approach is more natural and computationally straightforward when one is interested in low- to intermediate-resolution spectra. It also enables the treatment of more general initial states (e.g., thermal ensembles or non-Condon effects) and potential energy surfaces (e.g, Duschinsky rotation and anharmonicity).
@@ 2.2 Representation of SVL initial states by Hagedorn functions
in Hagedorn’s parametrization, where $x:=q-q_{t}$ is the shifted position, $q_{t}$ and $p_{t}$ are the position and momentum of the center of the wavepacket, $S_{t}$ is the classical action, and $Q_{t}$ and $P_{t}$ are complex-valued $D$-dimensional matrices that satisfy the so-called symplecticity conditions53, 46 (see eqs 5 and 6 of ref 43) and determine the width matrix $\mathrm{A}_{t}:=P_{t}\cdot Q_{t}^{-1}$ of the Gaussian. We can then apply Hagedorn’s raising operator
@@
in the form of multivariate polynomials multiplied by a common Gaussian, with $\langle j\rangle=(\underbrace{0,\dots,0}_{j-1},1,\underbrace{0,\dots,0}_{D-j})$ being the $D$-dimensional unit vector along the $j$-th degree of freedom.68, 44, 53, 46
The initial vibrational state $|\psi_{0}\rangle=|{K}\rangle$ in the SVL emission process can be exactly represented by a single Hagedorn function $\varphi_{K}$ in the normal-mode coordinates. In the harmonic approximation, normal-mode coordinates for a given electronic state diagonalize the Hessian at its equilibrium geometry, and the ground vibrational wavefunction is a Gaussian (eq 3) with a diagonal width matrix $\mathrm{A}_{0}=P_{0}\cdot Q_{0}^{-1}$. With both $Q_{0}$ and $P_{0}$ diagonal, the associated multidimensional Hagedorn function is a product of univariate functions, each of which is a Gaussian multiplied by a Hermite polynomial. As we will see in the next sections, this simple Hermite factorization is lost during the dynamics on a general harmonic surface.
@@ 2.3 Hagedorn wavepacket dynamics
Remarkably, Hagedorn wavepackets are exact solutions to the time-dependent Schrödinger equation (TDSE) with a harmonic potential
@@
$\dot{Q_{t}}$ $=m^{-1}\cdot P_{t},$ $\dot{P_{t}}$ $=-V_{g}^{\prime\prime}(q_{t})\cdot Q_{t}=-\kappa_{g}\cdot Q_{t},$
$\dot{S_{t}}$ $=L_{t}=p^{T}\cdot m^{-1}\cdot p/2-V_{g}(q_{t}),$ (7)
where $m$ is the mass matrix and $L_{t}$ is the Lagrangian.53, 46, 69 The multi-index $K$ of the Hagedorn function does not change during the exact propagation in up-to-quadratic potentials, i.e., the Hagedorn function propagated in a harmonic potential to time $t$ retains its form
$\varphi_{K,t}=(K!)^{-1/2}\left(A^{\dagger}[q_{t},p_{t},Q_{t},P_{t}]\right)^{K}\varphi_{0}[q_{t},p_{t},Q_{t},P_{t},S_{t}],$ (8)
even in the presence of mode mixing (when Hessian $\kappa_{g}$ is not a diagonal matrix). In fact, equations of motion (7) are the same as in the thawed Gaussian approximation,70, 71 which has been applied to vibronic spectra from ground vibrational levels23, 24, 72, 73 and extended to propagate not only Gaussians but also Gaussians multiplied by a linear polynomial prefactor.74, 30, 36, 32, 75, 76 Using Hagedorn wavepackets, it is possible to treat arbitrary polynomials times a Gaussian (describing, e.g., excited vibrational states), and since the propagation (7) still depends only on the Gaussian’s parameters, a single Gaussian trajectory is sufficient to obtain SVL spectra from all initial vibrational levels.
Although equations (7) can also be solved analytically to obtain a solution at an arbitrarily long time $t$ in a harmonic system,77 the spectrum (2) depends on the values of the autocorrelation function $C(t)$ at all times $t$ from 0 to $t_{\text{max}}$ (the time that gives the desired spectral resolution). Thus, compared to evaluating analytical solutions at each discretized time point over the total time of propagation, numerical propagation (with geometric integrators) is nearly as accurate and efficient. In high-dimensional harmonic systems, the computational cost of evaluating SVL spectra is dominated by computing the Hagedorn overlaps rather than the propagation, which only needs to be performed once for all initial states.
The numerical approach to wavepacket propagation becomes necessary when the Hagedorn dynamics is combined with the local harmonic approximation to take into account anharmonicity in flexible molecules. There, as in the thawed Gaussian approximation, only minimal modifications need to be made to eqs 7 by evaluating the gradient $V_{g}^{\prime}(q_{t})$ and Hessian $V_{g}^{\prime\prime}(q_{t})$ on the fly from electronic structure calculations (instead of from the static Hessian matrix $\kappa_{g}$ in the global harmonic model), but analytical solutions at arbitrary times are no longer possible. In the local harmonic case, the cost of evaluating SVL spectra is dominated by ab initio calculations and the cost of computing overlaps (in post-processing) becomes negligible.
Whereas closed-form analytical expressions exist for overlaps between two Gaussian wavepackets, computing the SVL spectra requires the overlaps between two “excited” Hagedorn functions associated with two different Gaussian centers, a problem similar to the computation of Franck–Condon factors in harmonic systems in the time-independent approach. In ref 61, we derived recurrence relations for algebraically evaluating the overlaps between any Hagedorn functions; these expressions are also presented succinctly in the Supporting Information here. Although more complex than simple Gaussian overlaps, these calculations typically involve only one or two initially excited modes containing just a few vibrational quanta.
Beyond the $O(D^{3})$ scaling with the number of normal modes $D$ (the same as in computing overlaps of Gaussian), the cost of evaluating the overlaps between Hagedorn functions depends on the number of excited modes and the vibrational excitation levels. For multiply excited levels in a single mode, the cost of evaluating $\langle\varphi_{K,t=0}|\varphi_{K,t}\rangle$ scales quadratically with the excitation $K$.
Compared to the sum-over-states approach, which requires overlaps between an initial vibrational excited state and a considerable number of possible final states in high-dimensional systems, the time-dependent Hagedorn method requires only overlaps between Hagedorn functions with the same multi-index $K$ (determined by the initial vibrational excitation). The recursive nature of the overlap expressions also means that computing autocorrelation functions for multiply excited levels inherently generates those for lower levels encountered during recursion (e.g., computing the autocorrelation for a doubly excited level also provides results for singly excited levels). Conversely, with efficient caching implemented, results for lower excitation levels can be reused for higher excitation levels.
@@ 2.4 Combination with ab initio electronic structure calculations
The first step in transforming Cartesian coordinates to mass-weighted normal-mode coordinates is minimizing the ro-vibrational coupling relative to a reference geometry. This is achieved by moving to the center-of-mass frame and applying the Kabsch algorithm to satisfy the translational and rotational Eckart conditions.78, 79, 80 Next, the normal modes and the corresponding transformation matrix are found by diagonalizing the mass-weighted Hessian matrix of the reference state and projecting out the rotational and translational degrees of freedom (see section 6.7 of ref 81). In the SVL application, the reference geometry and Hessian are taken from the initial, excited electronic state so that each vibrational eigenstate $|K\rangle$ may be represented by a single Hagedorn function $\varphi_{K}$. The diagonalized vibrational Hessian $\kappa_{e}$ in the excited-state normal-mode coordinates determines the width matrix $\mathrm{A_{0}}=i\sqrt{m^{1/2}\cdot\kappa_{e}\cdot m^{1/2}}$ of the Gaussian $\varphi_{0}$ associated with the SVL initial states. In Hagedorn’s parametrization, we set $Q_{0}=(\operatorname{Im}\mathrm{A}_{0})^{-1/2}$ and $P_{0}=\mathrm{A}_{0}\cdot Q_{0}$ to satisfy the symplecticity relations.
In the adiabatic harmonic approximation, the Hessian $\kappa_{g}$ defining the ground-state potential (eq 6) used for propagation is evaluated at the ground-state equillibrium geometry $q_{\text{ref},g}=q_{\text{eq},g}$; both the geometry and Hessian are transformed to be in the excited-state mass-weighted normal-mode coordinates. In the vertical harmonic approximation, the ground-state Hessian $\kappa_{g}$ is instead evaluated at the optimized excited-state geometry $q_{\text{ref},g}=q_{\text{eq},e}$. Due to mode mixing, $\kappa_{g}$ is in general not diagonal when expressed in the excited-state normal-mode coordinates. Therefore, the initially diagonal $Q_{t=0}$ matrix of the SVL initial state will no longer be diagonal after evolution with the ground-state potential, and the propagated Hagedorn function will not remain a simple product of univariate Hermite functions.47, 48, 46 However, a single Hagedorn function still suffices to represent the SVL process exactly in the harmonic limit. In contrast, basis expansion methods (using, e.g., products of Gaussians and Hermite polynomials) typically require many functions to maintain accuracy in the presence of mode mixing in multi-dimensional systems.
@@ 3 Computational details
To compare with previously reported results and to validate our method, we used the same density functional theory (DFT) method (at PBE0/def2-TZVP level of theory82, 83) as in ref 14 to construct the global harmonic models. Geometry optimizations and frequency calculations were performed using the Gaussian 16 package84 for the ground ($S_{0}\,\mathrm{{}^{1}A_{g}}$) and the first excited ($S_{1}\,\mathrm{{}^{1}B_{2u}}$) electronic states of anthracene. The excited-state calculations were carried out using standard linear-response time-dependent DFT. While this level of theory is somewhat simple and approximate, it was shown14 to reproduce emission spectra from the ground and singly excited levels reasonably well with results similar to those obtained from more expensive second-order approximate coupled-cluster calculations. The optimized structures of the two states and the frequencies of the vibrational modes analyzed in this work are listed in the Supporting Information. The assignment of the vibrational modes and their symmetry follows the convention of ref 3 and the Supporting Information of ref 14.
A second-order TVT geometric integrator 85, 53, 69 was used to propagate the parameters of the Gaussian for a total time of $8\times 10^{5}\,\text{au}$ ($\sim$19 ps) with a time step of 8 au in the global harmonic systems considered in this work. After the (single!) trajectory of the parameters of the Gaussian was obtained in a given harmonic system, the autocorrelation functions between Hagedorn wavepackets were computed every four steps for each initial vibrational level ($0^{0},12^{1},12^{2},\overline{11}^{1},\overline{11}^{2},12^{1}6^{1},12^{1}\overline{5}^{1}$, and $12^{1}5^{1}$) using the algebraic algorithm described in ref 61 (a Python implementation is provided in the supplementary material of ref 43). Given the limited range of the observed experimental data, it was unnecessary to extend the range of the simulated spectra by computing the autocorrelation function at every step. A Gaussian damping function with a half-width at half-maximum of 20000 au was applied to the autocorrelation functions before performing the Fourier transform.
To facilitate the comparison, the intensities of the highest peaks in all spectra (simulated and experimental) were set to unity. A horizontal shift, determined by aligning the $0^{0}_{0}$ peak in the computed ground-level ($0^{0}$) fluorescence spectrum (see the Supporting Information) to be at the experimental origin $27709\,\text{cm}^{-1}$, was applied to the wavenumbers of the simulated spectra in global harmonic potentials to correct for the relatively large error (compared to the vibrational features) in the computed electronic excitation energy; the $\omega^{3}$ factor in eq 2 was then adjusted based on the corrected wavenumbers (see the Supplemental Material of ref 62 for details).